products corporations have the following capital structure, which it considers optimal: Bonds, 7% (at par)                    Br 300,000   Preferred stock, Br.5                    240,000 Common stock                        360,000 Retained earnings                      300,000                                                   1,200,000 


1
Expert's answer
2022-01-02T18:20:27-0500

Let the dividends be 4 percent, and the price of stock is 40, tax rate 40%. Then:

"r=\frac{d}{P}=\frac{4}{40}=0.1"

Borrowed capital=300 000

Equity capital=240 000+360 000+300 000=900 000

Total capital=300 000+900 000=1 200 000

"WACC=10\times\frac{900 000}{1 200 000}+7\times{300 000}{1 200 000}\times(1-0.4)=8.55"